OpenStudy (anonymous):

Consider the system of equations. y = 2x + 2 y = 6x + 2 The graph of the system has ________ solutions. A. no B. exactly one C. infinite D. two

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

im thinking maybe B

OpenStudy (isaidavila):

Think is C

OpenStudy (anonymous):

if its b then do you have a pair

jhonyy9 (jhonyy9):

first you need solve the siystem and see what are the roots of

OpenStudy (welshfella):

HInt - put x = 0 what is the value of the y's?

OpenStudy (isaidavila):

Match both equation 6x+2=2x+2 6x=2x There is no solution

OpenStudy (anonymous):

I was directing @GalaxyKatz to that

OpenStudy (anonymous):

??

OpenStudy (anonymous):

So answer is a

OpenStudy (isaidavila):

For me yes

OpenStudy (welshfella):

No there is one solution (0,2)

OpenStudy (welshfella):

draw the graphs and you will see they intersect at (0,2)

OpenStudy (anonymous):

^^^ i think you may be right @welshfella

OpenStudy (welshfella):

6x + 2 is equal to 2x + 2 when x = 0

OpenStudy (isaidavila):

Oh thats true @welshfella

OpenStudy (anonymous):

thanks guys!!

OpenStudy (welshfella):

if we subtract the 2 equations we get 4x = 0 so x = 0

OpenStudy (welshfella):

yw